# RGBDS-SLAM: A RGB-D Semantic Dense SLAM Based on 3D Multi Level Pyramid Gaussian Splatting

### v1

@@ Abstract
High-quality reconstruction is crucial for dense SLAM. Recent popular approaches utilize 3D Gaussian Splatting (3D GS) techniques for RGB, depth, and semantic reconstruction of scenes. However, these methods often overlook issues of detail and consistency in different parts of the scene. To address this, we propose RGBDS-SLAM, a RGB-D semantic dense SLAM system based on 3D multi-level pyramid gaussian splatting, which enables high-quality dense reconstruction of scene RGB, depth, and semantics.In this system, we introduce a 3D multi-level pyramid gaussian splatting method that restores scene details by extracting multi-level image pyramids for gaussian splatting training, ensuring consistency in RGB, depth, and semantic reconstructions. Additionally, we design a tightly-coupled multi-features reconstruction optimization mechanism, allowing the reconstruction accuracy of RGB, depth, and semantic maps to mutually enhance each other during the rendering optimization process. Extensive quantitative, qualitative, and ablation experiments on the Replica and ScanNet public datasets demonstrate that our proposed method outperforms current state-of-the-art methods. The open-source code will be available at: https://github.com/zhenzhongcao/RGBDS-SLAM.
@@ II-A NeRF-based SLAM
The development in neural implicit representations, particularly those based on NeRF, have significantly enhanced the performance of SLAM systems. Among them, NICE-SLAM[8] is the first solution to combine NeRF and SLAM, which incorporates multi-level local information by introducing a hierarchical scene representation, enabling efficient map construction and robust tracking. However, NICE-SLAM suffers from computational efficiency issues. Therefore, [9], [10], [11] and [12] have introduced voxel-based neural representations, coordinate and sparse parameters, hybrid representation of signed distance fields (SDF) and neural point cloud respectively to optimize and improve computational efficiency. The above solutions do not consider semantic mapping, so based on these solutions, NIDS-SLAM[13] introduce a novel approach for dense 3D semantic segmentation, based on 2D semantic color maps of keyframes, which are able to accurately learn the dense 3D semantics of the scene online while simultaneously learning geometry. However, this work does not integrate semantic with other features of the environment, such as geometry and appearance. Therefore, DNS-SLAM[14] integrates multi-view geometry constraints with image-based feature extraction to improve appearance details and to output color, density, and semantic class information. SNI-SLAM[15] introduce cross-attention based feature fusion to incorporate semantic, appearance, and geometry features, thus improving the accuracy of mapping, tracking, and semantic segmentaion. Although these NeRF-based SLAM schemes achieve high-quality reconstruction effects, they suffer from poor scalability, low efficiency and poor real-time performance due to NeRF.
@@ III-A Overall System Framework
Therefore, we construct an n-layer image pyramid for RGB, depth, and semantic maps.
@@ V Conclusion
In this paper, we propose RGBDS-SLAM, which is a complete RGB-D semantic dense SLAM system. We introduce a 3D multi-level pyramid gaussian splatting method to reconstruct the details and consistency of the scene. We also design a tightly coupled multi-feature reconstruction optimization mechanism that promotes the optimization of RGB, depth, and semantic features, enhancing each other. Experiments also demonstrate the effectiveness and scalability of our proposed method. However, we have not considered the issue of dynamic scenes. Robustly reconstructing the RGB, depth, and semantic information in dynamic scenes will be the focus of our future work.

### v2

@@ Abstract
High-fidelity reconstruction is crucial for dense SLAM. Recent popular methods utilize 3D gaussian splatting (3D GS) techniques for RGB, depth, and semantic reconstruction of scenes. However, these methods ignore issues of detail and consistency in different parts of the scene. To address this, we propose RGBDS-SLAM, a RGB-D semantic dense SLAM system based on 3D multi-level pyramid gaussian splatting, which enables high-fidelity dense reconstruction of scene RGB, depth, and semantics. In this system, we introduce a 3D multi-level pyramid gaussian splatting method that restores scene details by extracting multi-level image pyramids for gaussian splatting training, ensuring consistency in RGB, depth, and semantic reconstructions. Additionally, we design a tightly-coupled multi-features reconstruction optimization mechanism, allowing the reconstruction accuracy of RGB, depth, and semantic features to mutually enhance each other during the rendering optimization process. Extensive quantitative, qualitative, and ablation experiments on the Replica and ScanNet public datasets demonstrate that our proposed method outperforms current state-of-the-art methods, which achieves great improvement by 11.13% in PSNR and 68.57% in LPIPS. The open-source code will be available at: https://github.com/zhenzhongcao/RGBDS-SLAM.
@@ II-A NeRF-based SLAM
The development in neural implicit representations, particularly those based on NeRF, have significantly enhanced the performance of SLAM systems. Among them, NICE-SLAM[8] is the first solution to combine NeRF and SLAM, which incorporates multi-level local information by introducing a hierarchical scene representation, enabling efficient map construction and robust tracking. However, NICE-SLAM suffers from computational efficiency issues. Therefore, [9], [10], [11] and [12] have introduced voxel-based neural representations, coordinate and sparse parameters, hybrid representation of signed distance fields (SDF) and neural point cloud respectively to optimize and improve computational efficiency. The above solutions do not consider semantic mapping, so based on these solutions, NIDS-SLAM[13] introduce a novel approach for dense 3D semantic segmentation, based on 2D semantic color information of keyframes, which are able to accurately learn the dense 3D semantics of the scene online while simultaneously learning geometry. However, this work does not integrate semantic with other features of the environment, such as geometry and appearance. Therefore, DNS-SLAM[14] integrates multi-view geometry constraints with image-based feature extraction to improve appearance details and to output color, density, and semantic class information. SNI-SLAM[15] introduce cross-attention based feature fusion to incorporate semantic, appearance, and geometry features, thus improving the accuracy of mapping, tracking, and semantic segmentaion. Although these NeRF-based SLAM schemes achieve high-quality reconstruction effects, they suffer from poor scalability, low efficiency and poor real-time performance due to NeRF.
@@ III-A Overall System Framework
Therefore, we construct an n-layer image pyramid for RGB, depth, and semantic images.
@@ V Conclusion
In this paper, we propose RGBDS-SLAM, which is a complete RGB-D semantic dense SLAM system, focusing on gaussian mapping. We first introduce a 3D multi-level pyramid gaussian splatting method to reconstruct the details and consistency of the scene. We futhermore design a tightly coupled multi-feature reconstruction optimization mechanism that promotes the optimization of RGB, depth, and semantic features, enhancing each other. Experiments also demonstrate the effectiveness and scalability of our proposed method. However, we have not considered the issue of dynamic scenes. Robustly reconstructing the RGB, depth, and semantic information in dynamic scenes will be the focus of our future work.